OpenStudy (anonymous):

Find the derivative of: f(x)=sin(x)+cos(x) Find the derivative of: f(x)=sin(x)+cos(x) @Mathematics

OpenStudy (anonymous):

Is it f '(x)=0?

jimthompson5910 (jim_thompson5910):

f(x)=sin(x)+cos(x) f'(x)=d/dx(sin(x)+cos(x)) f'(x)=d/dx(sin(x))+d/dx(cos(x)) f'(x)=cos(x) - sin(x)

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

Thanks